HUBBLE SPACE TELESCOPE DAILY REPORT #2261 PERIOD COVERED: 0000Z (UTC) 12/01/98 - 0000Z (UTC) 12/02/98 Daily Status Report as of 336/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Fourteen Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.3 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 6745 (Fundamental Plane, Morphology-Density Relation, and Lensing in the z=0.58 Arc Cluster CL2053) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-2053 in order to study its fundamental plane structure, its morphology-density relation, and its lensing. The distribution of {post}starburst galaxies and/or mergers as a function of radius will be of particular interest. The mass distribution will be determined from the distortion of background galaxies by weak lensing. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Eight Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.6 Completed STIS/CCD 7930 (STIS CCD Binned Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make numerous bias exposures in 1 x 2 binned mode in order to permit construction of "superbias" frames in which the effects of read noise have been rendered negligible. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Five Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.9 Completed STIS/CCD 7310 (The Interaction of Accretion Disks with Protostellar Binaries) The Space Telescope Imaging Spectrograph (CCD) was used to study the interaction of the accretion disk with the protostellar binary system 040142+215. This should determine how the accretion disk interacts with stars in young binary systems and will measure the rates that mass accumulates onto each component of the pair. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 6754 (Imaging of the HH 30 Circumstellar Disk and Jets) The WF/PC-2 was used to make observations of the disk and jets of the Herbig-Harrow Object HH-30. Recent Hubble Space Telescope images have revealed HH-30 to be a prototype young stellar object system possessing an optically thick circumstellar absorption disk extended perpendicular to highly collimated bipolar jets. The observations were executed as scheduled, and no anomalies were noted. HH-30 1.11 Completed WF/PC-2 6347 (A Search for Jets in Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebulae (class PNe) NGC-2392 in order to search for the presence of possible narrow jets. The current model for PNe formation indicates that jets are an unexpected phenomena, yet there is a growing body of circumstantial evidence that suggests narrow jets are common feature in PNe. The observations were completed as scheduled, and no anomalies were noted. 1.12 Completed STIS/CCD/MA1/MA2 7269 (The 1997/98 Eclipse of VV Cephei) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make ultraviolet spectrographic observations of the long-period eclipsing binary star HD208816 (a.k.a., VV-Cephei). This eclipse will provide an unprecedented opportunity to obtain spatially resolved information about the density, ionization and velocity structure in the outer atmosphere of the M supergiant primary. A wave length calibration observation was also taken. The observations were completed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.14 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: The SSR EDAC error counter was cleared four times per ROP SR-1. Using ROP IC-2. the SSA transmitter was turned off and/or on twice. Two TTRs were written to document a total of seven missing minor frames. As described in HSTAR 6794, at 336/064955Z there was a NICMOS single bit error. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH